# ©Plug-in Authorization for Human Content Copyright Protection in Text-to-Image Model

### v1

@@ 2.2 Extraction: Reverse LoRA
Overall, after the de-concept step, the model $w-{w}_{L}$ is unable to generate images in the Picasso style, and it performs well with surrounding prompts thanks to the re-context step. Therefore, through the extraction operation, we obtain a non-infringing model $\tilde{w}=w-{w}_{L}$ and a ©plug-in ${w}_{L}$. By incorporating the ©plug-in, the model is restored to the original base model $w$, regaining the capability of successfully generating the artworks in the “Picasso” style. The intermediate results of the extraction process are visually illustrated in Figure 5 in Appendix B, showcasing the successful extraction of the targeted copyright, while preserving the model’s ability to generate images with surrounding contexts.
@@ 2.3 Combination: EasyMerge
$\mathop{\arg\min}\limits_{w_{L}}\ \sum_{c_{k}\in S,j\in S_{L}}\mathbb{E}_{\epsilon,t}\|{\phi^{j}_{(w+w_{L})}(\epsilon,c_{k},t)-\phi^{j}_{(w+w_{L_{k}})}(\epsilon,c_{k},t)}\|^{2},$ (5)
where $\phi^{j}$ is the output of layer $j$’s LoRA component, $w_{L_{k}}$ is the LoRA plug-in of the context $c_{k}$, ${w}_{L}$ denotes the combined plug-in, $S$ is the set of target concepts $c_{k}$ to be combined, $S_{L}$ is the set of layers that are equipped with LoRA components. Similarly to the previous section, $w$ denotes the base model parameter, $\epsilon$ is initial noise and $t$ is the sampling timestep of the diffusion process. The implementation details are described in Algorithm 1.
@@
0:  A set $S$ of indices of plug-ins to be combined, a set $S_{L}$ of indices of layers with LoRA components, a base model $w$ and a diffusion step $T$
0:  Combined LoRA plug-in $w_{L}$
1:  while not converge do
2:     for $c_{k}\in S$ do
3:        ## The pair of plug-in and prompt to be combined, i.e., $(w_{L_{k}},c_{k})$
4:        Sample $t\sim\text{Uniform}([1...T]);\epsilon\sim N(\mathbf{0},\mathbf{I})$
5:        Record the input and output of each layer $j$: $I_{w_{L_{k}}}^{j},O_{w_{L_{k}}}^{j}\leftarrow\phi^{j}_{(w+w_{L_{k}})}(\epsilon,c_{k},t)$
6:        Compute the output of $w_{L}$: $O_{w_{L}}^{j}\leftarrow\phi_{{w_{L}}}^{j}(I_{w_{L_{k}}}^{j})$
7:        Compute the loss: $\mathcal{L}\leftarrow\sum_{j\in S_{L}}\|O_{w_{L}}^{j}-O_{w_{L_{k}}}^{j}\|^{2}$
8:        Update the parameter: $w_{L}\leftarrow w_{L}-\nabla_{w_{L}}\mathcal{L}$
9:     end for
10:  end while
We note that the above combination process can also be applied to the scenario of extracting multiple concepts simultaneously. The non-infringing model $w-w_{L}$ represents a model that simultaneously excludes multiple contexts $c_{k}$ given each extraction plugin $w_{L_{k}}$, which is called the combination of extraction.
3 Results
@@
Given that the addition operation has already been well demonstrated by existing practices, we focus on evaluating the extraction and combination operations. We choose two typical scenarios of copyright infringement: artist-style replication and cartoon intellectual property (IP) recreation.
3.1 Experiment Setup, Metrics and Baseline Comparison
Experiment Setup. Our experiments involve fine-tuning the U-Net architecture of the Stable Diffusion Model v1.5. For extraction, we generate data with pre-trained models, using ChatGPT to create content related to both artistic styles and specific IP characters. We conduct a 10-epoch training with 30 iterations per epoch, employing a learning rate of 1.5e-4 and a rank of 40 for LoRA. The combination operation uses a learning rate of 1e-3 and a rank value of 140 for LoRA.
Evaluation Metrics. The effectiveness of extraction is measured by the discrepancy between images generated by the base model and the non-infringing model post-extraction, using the same prompts. We employ Kernel Inception Distance (KID) [23] and Learned Perceptual Image Patch Similarity (LPIPS) [24] as objective metrics, alongside human judgment, to evaluate the quality and stylistic consistency of generated images.
@@ 3.1 Experiment Setup, Metrics and Baseline Comparison
Baseline Comparison. We compare our extraction operation with the concept ablation approach [25] and Erased Stable Diffusion (ESD) [26], which achieve concept removal by aligning latent representations of target concepts with those of anchor concepts. Due to the complexity of image generation setups, we conservatively compare our methods with them by using similar scenarios and metrics.
3.2 Extraction and Combination of Artists’ Styles Replication
@@
Figure 3: Results of style replication. In Figure (a), we show samples after extracting the different artist’s style in each column. Each non-infringing model exhibits a deficiency in one style generation ability, with all other style generation capabilities remaining unaffected. In Figure (b), we present samples generated after using certain ©Plug-ins in each column. Each of these ©Plug-ins serves to exclusively restore the generation of one particular style, while the generation of other styles continues to exhibit diminished performance.
@@
Extract artist styles from the Stable Diffusion V1.5 (base model). We consider three renowned artists: (1) Vincent van Gogh, (2) Pablo Ruiz Picasso and (3) Oscar-Claude Monet. The results of individual style extraction are visually presented in Figure 3(a), where we compare images generated by the base model and the non-infringing model using target and surrounding style prompts. The images enclosed in red boxes represent the target styles, showcasing a marked contrast with the base model’s outputs. This contrast confirms the successful isolation of the target styles. Meanwhile, images for surrounding styles maintain a high degree of similarity with the base model, indicating the preservation of image quality for non-target styles.
@@
Quantitative analysis, presented in Table 1, further validates the effectiveness of our extraction operation. The Kernel Inception Distance (KID) metric [23] quantifies the discrepancy between the set of images generated by the base model and that generated by the non-infringing model after extraction with the same set of prompts. KID is akin to the Fréchet Inception Distance (FID) [27] but is considered to exhibit less bias. The KID of our extraction method significantly outperforms the Concepts-Ablation method [25] by increasing from 42 to 187 for the target style, indicating more effective removal of target concepts. Additionally, our method reduces the Learned Perceptual Image Patch Similarity (LPIPS) from 0.21 to 0.14 for surrounding styles in comparison to the Erasing Stable Diffusion (ESD) method [26], indicating less degradation and a better preservation of the artistic quality. LPIPS is a robust measurement tool that effectively captures differences in human perception between two images, offering a comprehensive evaluation of stylistic variations in generated artworks.
@@
Combination of extractions. We sequentially extract the styles of Van Gogh, Picasso, and Monet, and build a non-infringing model by using combination of extractions. Then we reintroduce them by using their corresponding copyright plug-ins. The non-infringing model, shown in the leftmost column of Figure 3(b), generates images that are notably different from those of the base model, highlighting the effectiveness of the multiple extraction operations.
Upon using the individual style copyright plug-ins, the images within the red boxes exhibit the target styles with a striking resemblance to the base model’s outputs. This outcome demonstrates that ©plug-ins can restore the model’s capability to generate artwork in the target styles while adhering to copyright restrictions.
3.3 Extraction and Combination of Cartoon IPs Recreation
In the realm of intellectual property (IP) recreation, extraction and combination operations showcase significant capabilities. Figure 4(a) presents the extraction results for three iconic IP characters: Mickey Mouse, R2D2, and Snoopy. The red-boxed images, generated after extracting the target IP prompts, deviate markedly from the base model’s outputs, demonstrating the effectiveness of target extraction. Notably, other IP images maintain similarity with the base model, confirming that the extraction is precise and does not disrupt the generation of surronding IP-related content.
Quantitative analysis in Table 2 reveals a substantial 2.6-fold increase in the Kernel Inception Distance (KID) score for the targeted IPs, while other IPs’ KID metrics remain stable, underscoring the method’s targeted isolation of specified IPs.
@@ 3.3 Extraction and Combination of Cartoon IPs Recreation
Figure 4: Results of IP Recreation. In Figure (a), each column on the right represents the output after extraction. We successfully extract the unique IPs of Mickey, R2D2, and Vader independently, preserving the generation of other IPs. In Figure(b), we generate either Mickey or Vader in a single image by using the corresponding Plug-in and generate them both by using the combined ©plug-in.
Moreover, our extensive large-scale experiments confirm that the extraction process does not affect the model’s capacity to produce content unrelated to particular IPs. The model’s performance post-extraction remains stable across routine or common usage scenarios, as elaborated in the Appendix C. This guarantees that while the model adheres to copyright limitations by appropriately handling or isolating particular IPs, it maintains its overall usefulness and diverse creative capabilities.
@@
The combination operation is illustrated in Figure 4(b), where the initial image lacks recognizable IPs. The subsequent images, using individual and combined copyright plug-ins for Mickey and Vader, clearly display the respective IPs. The final image is generated upon using the combined ©plug-in, successfully displaying both IPs. This sequence highlights the plug-in’s efficacy in selectively and collectively restoring the model’s ability to generate IP-specific content.
These results indicate the efficacy of combining multiple extractions, where the non-infringing model’s ability to produce images themed with either Mickey Mouse or Darth Vader is disabled. With the use of the respective copyright plug-in, the model regains the capability to create content related to the specific IP.
4 Discussion
The rapid advancement of generative AI models has sparked significant concern over their potential to produce copyright-infringing content. As these models elevate the quality of their outputs, the lack of proper attribution to human creators becomes increasingly problematic. To mitigate these societal concerns, we introduce the “Copyright Plug-in Authorization” framework, a system designed to align with the principles of copyright law and ensure that the use of copyrighted material in AI-generated content is appropriately managed and rewarded.
One instantiation of our framework leverages the Low-Rank Adaptor (LoRA) to convert copyrighted data into plug-ins, facilitating the tracking of content usage and the fair distribution of rewards to original creators. This approach not only respects the rights of content creators but also provides a clear and accessible mechanism for users to engage with copyrighted material in a legally compliant manner.
Nonetheless, the current framework may still face challenges when managing a vast array of plug-ins. As the number of plug-ins grows, maintaining a user-friendly interface and ensuring seamless access to specific generations becomes critical. This requires innovative solutions to organize and retrieve plug-ins effectively, without compromising the user experience. Another challenge is ensuring the backward compatibility of plug-ins when the base model is updated. Retraining the entire collection of plug-ins following a model update can be prohibitively expensive and time-consuming. Ensuring backward compatibility, so that previously created plug-ins continue to function with new model iterations, is essential to avoid obsolescence and maintain user trust.
Our Copyright Plug-in Authorization framework represents an initial step towards balancing the innovative potential of generative AI with the imperative of copyright protection. By addressing the challenges of plug-in management, model updating, and performance maintenance, we aim to foster an ecosystem where creativity and legal compliance can coexist and thrive.
Appendix A Related Work
In order to position our work in the vast literature, we review related work through two perspectives: scope and technique. It is worthy to note that some of the literature touch both sides and we organize them in a way most related to ours.
A.1 Scope Related: Copyright, Data Contribution and Credit Attribution
Recent text-to-image generative models are trained with large scale datasets [1, 28], which cannot be guaranteed free of copyrighted data. At the same time, the state-of-the-art models are capable of generating high-quality and valuable creative images comparable to human creators or even memorizing the data points in the training set [29], which arouses copyright concerns about the training data and brings anxiety to the artist community.
@@ A.1 Scope Related: Copyright, Data Contribution and Credit Attribution
Shan et al. [35] propose Image Cloaking that suggests adding adversarial perturbations before posting artistic works on the internet so as to make them unlearnable for machine learning model, which has been pointed out to be hard to defend against future learning algorithms [36].
@@
Theoretically, [37, 38] connect the copyright protection of training data with the concept of differential privacy and discuss their subtle differences. [39] further formulate the copyright problem with a near free access (NAF) notion to bound the distance of the generative distributions of the models trained with and without the copyrighted data.
@@
Our paper distinguishes largely from all previous works as we do not try to prohibit generating copyrighted concepts but instead we introduce a copyright authorization for the generative model to reward the copyright owners with fairness and transparency. From this aspect, our paper is also related with literature of monetizing the training data [40, 41, 42, 43] or attributing credits for the generative contents [8], but we establish a very distinct way to reward the authorship.
@@
Heated discussion is also around the copyright for AI generated art work Franceschelli and Musolesi [44], Abbott and Rothman [45]. The Review Board of the United States Copyright Office has recently refused the copyright registration of a two-dimensional AI generated artwork entitled “A Recent Entrance to Paradise”. However, [45] argues for giving the copyright to AI generated works, which will encourage people to develop and use creative AI, promote transparency and eventually benefit the public interest.
A.2 Technique Related: Concept Removal, and Negative Sampling

### v2

@@ 2.2.2 Step2: Re-context
Overall, after the de-concept step, the model $w-{w}_{L}$ is unable to generate images in the Picasso style, yet it performs well with surrounding prompts thanks to the re-context step. Therefore, through the extraction operation, we obtain a non-infringing model $\tilde{w}=w-{w}_{L}$ and a ©plug-in ${w}_{L}$. By incorporating the ©plug-in, the model is restored to the original base model $w$, regaining the capability of successfully generating the artworks in the “Picasso” style. The intermediate results of the extraction process are visually illustrated in Figure 7 in Appendix A.1, showcasing the successful extraction of the targeted copyright, while preserving the model’s ability to generate images with surrounding contexts.
@@ 2.3 Combination: EasyMerge
$\mathop{\arg\min}\limits_{w_{L}}\ \sum_{k\in S,j\in S_{L}}\mathbb{E}_{\epsilon,t}\|{\phi^{j}_{w-w_{L}}(\epsilon,c_{k},t)-\phi^{j}_{w-w_{L_{k}}}(\epsilon,c_{k},t)}\|^{2},$ (5)
where $S$ is the set of text prompts to be combined, $S_{L}$ is the set of layers that are added with LoRA components, and $\phi^{j}$ is the output of layer $j$’s LoRA component. Similarly to the previous section, $w$ denotes the base model parameter, ${w}_{L}$ denotes the combined plug-in, $c_{k}$ denotes the prompt $k$, $w_{L_{k}}$ is the plug-in of context $c_{k}$, $\epsilon$ is initial noise and $t$ is the sampling timestep of the diffusion process. The non-infringing model $w-w_{L}$ is that simultaneously excludes multiple styles related to $c_{k}$, which is called the combination of extraction. Algorithm 1 describes concrete steps of optimizing the objective (5).
Input: A set $S$ of indices of plug-ins to be combined, base model $w$, diffusion step $T$
Output: Combined LoRA $w_{L}$
repeat
       for $w_{L_{i}},c_{i}\in S$ do
             $t\sim\text{Uniform}([1...T])$;
             $\epsilon\sim\mathcal{N}(0,1)$;
             AddHook($w_{L_{i}}$) ;
              // Capture input $I_{w_{L_{i}}}^{j}$ and output $O_{w_{L_{i}}}^{j}$ for each layer $j$
             $I_{w_{L_{i}}}^{j},O_{w_{L_{i}}}^{j}\leftarrow\Phi_{w+w_{L_{i}}}(\epsilon,c_{i},t)$ ;
              // Denoise to obtain features
             $O_{w_{L}}^{j}\leftarrow\phi_{{w_{L}}}^{j}(I_{w_{L_{i}}}^{j})$ ;
              // Get layer-output through new LoRA
             $\mathcal{L}\leftarrow\sum_{j\in S_{L}}\|O_{w_{L}}^{j}-O_{w_{L_{i}}}^{j}\|$;
             $w_{L}\leftarrow w_{L}-\nabla_{w_{L}}\mathcal{L}$;
       end for
until convergence;
@@
3 Experiments to Verify Basic Operations
@@
As a position paper, we regard our primary contribution as the proposal of the copyright authorization framework. Nonetheless, we also aim to validate the practical effectiveness and efficiency of the proposed basic operations. Given that the addition operation has already been well demonstrated by existing practices, we focus on evaluating the extraction and combination operations. We choose two typical scenarios of copyright infringement: artist-style replication and cartoon intellectual property (IP) recreation.
3.1 Experiment Setup, Metrics and Baselines
Experiment Setup. In all experiments, we fine-tune the attention component in the U-Net architecture of Stable Diffusion Model v1.5, as described in Rombach et al. (2022a).
For the extraction operation, we need to generate data with pre-trained models. For the case of extracting a given artistic style, we leverage ChatGPT (cha, ) to generate 10 common content in paintings. In the de-concept step, for each iteration, we select one of these content to generate 8 images with prompts “ The painting of [content] by [artist]”. Similarly, during re-context step, we select imagery to generate 8 images with prompts “ The painting of [content]” while using negative prompts “by [artist]”. For the case of extracting a particular IP character, it follows the same procedure as above except that the prompts become “The cartoon of the [IP character]” for the de-concept process and “The cartoon of the [character]” for the re-context process, respectively. For both the de-concept process and the re-context process, the training consists of 10 iterations, with each iteration 30 epochs. We use a learning rate of 1.5e-4, $T=50$ steps for the diffusion process, and a rank of 40 for LoRA.
For the combination operation, we use a learning rate of 1e-3 and a rank value of 140 for LoRA.
Metric. To evaluate the effectiveness of the extraction operation, we measure the discrepancy between the set of images generated by the base model and that generated by the non-infringing model after extraction with the same set of prompts. We want to observe a large discrepancy when the prompts are with target concepts while having a small discrepancy with surrounding concepts. This means that the extraction operation achieves its goal: the non-infringing model cannot generate images with target concepts but can generate high-quality images with surrounding prompts.
We acknowledge that for image generation tasks, the ultimate evaluation criterion is human judgment. Therefore, we provide the generated images from various scenarios for readers’ assessment. Nevertheless, to reduce costs and facilitate comparisons with existing approaches, we also employ an objective metric known as the Kernel Inception Distance (KID) (Bińkowski et al., 2018) to quantify the aforementioned discrepancy. KID is akin to the Fréchet Inception Distance (FID) (Heusel et al., 2017) but is considered to exhibit less bias and possess asymptotical normality. Moreover, we also employ the Learned Perceptual Image Patch Similarity (LPIPS) (Zhang et al., 2018) to quantify the discrepancy of artistic style artworks. LPIPS is a robust measurement tool that effectively captures differences in human perception between two images, offering a comprehensive evaluation of stylistic variations in generated artworks.
@@ 3.1 Experiment Setup, Metrics and Baselines
Baseline. We compare our extraction operation with the concept ablation approach (Kumari et al., 2023) and Erased Stable Diffusion (ESD) (Rohit Gandikota and Joanna Materzyńska and Jaden Fiotto-Kaufman and David Bau, 2023), which achieve concept removal by aligning latent representations of target concepts with those of anchor concepts.
In general, we find it hard to compare the results with existing methods because of the complex setups in image generation, e.g., the tuning steps and the trade-off between removing the target concept and keeping the surrounding concept. Therefore we take a conservative approach and only consider the generation with similar scenarios and the same metric as in the original paper.
3.2 Extraction and Combination of Artists’ Styles
(a) Results of extraction in style replication
(b) Results of combination in style replication
@@ 3.2 Extraction and Combination of Artists’ Styles
Figure 4: Results of style replication. In Figure (a), we show samples from different non-infringing models in each column. Each non-infringing model exhibits a deficiency in one style generation ability, with all other style generation capabilities remaining unaffected. In Figure (b), we present samples generated after integrating certain ©Plug-ins in each column. Each of these ©Plug-ins serves to exclusively restore the generation of one particular style, while the generation of other styles continues to exhibit diminished performance.
@@
Extraction. We extract artist styles from the Stable Diffusion V1.5, referred to as the “base model”. We consider three renowned artists: (1) Vincent van Gogh, (2) Pablo Ruiz Picasso and (3) Oscar-Claude Monet. The results of individual style extractions are visually presented in Figure 4(a). These images showcase outputs generated by both the base model and the non-infringing model, encompassing both the target styles and surrounding styles.
In Figure 4, images within red boxes represent the target styles, while the rest images embody surrounding styles. A notable contrast is observable between the images within the red boxes and those generated by the base model. However, the images representing surrounding styles exhibit a substantial similarity to those generated by the base model. This demonstrates the success of the extraction operation in isolating the target style from the base model while preserving the quality of images with surrounding styles.
@@
In Table 1, we employ quantitative metrics to evaluate the effectiveness of the extraction operation in comparison to baseline methods. Our method demonstrates notable improvements, as indicated by the KID metric increasing from 42 to 187 for the target style when compared to Concepts-Ablation (Kumari et al., 2023). This increase indicates an enhanced removal of the target style. Additionally, in a comparative assessment with the Erasing method (Rohit Gandikota and Joanna Materzyńska and Jaden Fiotto-Kaufman and David Bau, 2023), our method achieves a reduction in LPIPS from 0.21 to 0.14 for surrounding styles. This reduction implies less degradation of the surrounding artistic styles, affirming our method’s ability to preserve the quality of generated images when using surrounding style prompts.
Combination. In this part, we show the effectiveness of the combination operation for extracting multiple artist styles and then adding them back with corresponding plug-ins.
@@
Given three artistic styles of Van Gogh, Picasso, and Monet and their ©plug-ins, we first extract these three styles from the base model, producing a non-infringing model, which is illustrated in the leftmost column of Figure 4(b). Notably, all the images generated by the non-infringing model significantly differ from those generated by the base model in the leftmost column of Figure 4(a). This underscores the efficacy of the combination of multiple extraction operations.
We then individually integrate each style copyright plug-in into the non-infringing model. The images highlighted within red boxes represent the target style achieved after integrating the respective copyright plug-in. Notably, the target style images after integration are distinctly different from those produced by the non-infringing model, showing a closer resemblance to the images generated by the base model. This observation indicates that the copyright plug-in can reinstate the model’s ability to create artworks in the target styles, without infringing upon the copyright restrictions associated with other artistic styles.
3.3 Extraction and Combination of Cartoon IPs
In the context of intellectual property (IP) recreation, we demonstrate the capabilities of our framework through both extraction and combination operations. Specifically, Figure 5 displays the outcomes of extracting three iconic IP characters: Mickey, R2D2, and Snoopy. The images framed in red boxes were generated by the non-infringing model using prompts specific to the target IP, after extraction. These images significantly deviate from those produced by the base model. In contrast, images outside the red boxes, which represent other IPs, show a resemblance to those generated by the base model, indicating the targeted nature of the extraction process.
@@ 3.3 Extraction and Combination of Cartoon IPs
Figure 5: Results of IP Recreation. Each column on the right represents the output of a distinct non-infringing model. We successfully extract the unique IPs of Mickey, R2D2, and Vader independently, preserving the generation of other IPs.
Our approach to IP extraction effectively isolates the specified IP, ensuring that the generation capabilities for other IPs remain intact. The efficacy of our extraction method in the realm of IP recreation is quantified in Table 2, where we document a notable improvement, i.e., approximately a 2.6-fold increase in the Kernel Inception Distance (KID) metric for the targeted IP, while the KID metrics for other IPs remain relatively stable.
Additionally, our comprehensive large-scale experiments demonstrate that the extraction process does not impact the model’s ability to generate non-IP-related content. The performance of the model after extraction remains consistent with routine or everyday use cases, as further detailed in the appendix A.2. This ensures that while the model respects copyright constraints by effectively removing or isolating specific IPs, it does not compromise on its general utility or the breadth of its creative outputs.
@@
Furthermore, we demonstrate the combination of multiple IP ©plug-ins, as illustrated in Figure 6. The initial image is produced by the non-infringing model after extracting Mickey Mouse and Darth Vader, where the IPs are hard to recognize. Subsequent images, the second and third, are created after individually adding the Mickey and Vader ©plug-ins, which distinctly feature the respective IP. The final image is generated upon adding the combined ©plug-in, successfully displaying both IPs. This procedure underscores the plug-in’s efficacy in selectively and collectively restoring the model’s ability to generate IP-specific content.
Figure 6: IP Recreation in a single image. We can integrate ©plug-in into the non-infringing model to generate either Mickey or Vader in a single image or integrate the combined ©plug-in to generate both of them.
These results indicate the efficacy of combining multiple extractions, where the non-infringing model’s ability to produce images themed with either Mickey Mouse or Darth Vader is disabled. With the integration of the respective copyright plug-in, the model regains the capability to create content related to the specific IP.
4 Related Work
To position our work in the vast literature, we review related work through two perspectives: scope and technique. It is worth noting that some of the literature touches both sides and we organize them in a way most related to ours.
4.1 Scope Related: Copyright, Data Contribution and Credit Attribution
Recent text-to-image generative models are trained with large-scale datasets (Schuhmann et al., 2022; Liu et al., 2022), which cannot be guaranteed free of copyrighted data. At the same time, the state-of-the-art models are capable of generating high-quality and valuable creative images comparable to human creators or even memorizing the data points in the training set (Carlini et al., 2023b), which arouses copyright concerns about the training data and brings anxiety to the artist community.
@@ 4.1 Scope Related: Copyright, Data Contribution and Credit Attribution
Shan et al. (2023) propose Image Cloaking that suggests adding adversarial perturbations before posting artistic works on the internet so as to make them unlearnable for machine learning model, which has been pointed out to be hard to defend against future learning algorithms (Radiya-Dixit et al., 2021).
@@
Theoretically, Bousquet et al. (2020); Elkin-Koren et al. (2023) connect the copyright protection of training data with the concept of differential privacy and discuss their subtle differences. Vyas et al. (2023) further formulate the copyright problem with a near free access (NAF) notion to bound the distance of the generative distributions of the models trained with and without the copyrighted data.
@@
Our paper distinguishes largely from all previous works as we do not try to prohibit generating copyrighted concepts but instead we introduce a copyright authorization for the generative model to reward the copyright owners with fairness and transparency. From this aspect, our paper is also related with literature of monetizing the training data (Vincent & Hecht, 2021; Vincent et al., 2021; Li et al., 2022b; a) or attributing credits for the generative contents (Troynikov, 2023), but we establish a very distinct way to reward the authorship.
@@
Heated discussion is also around the copyright for AI generated artwork Franceschelli & Musolesi (2022); Abbott & Rothman (2022). The Review Board of the United States Copyright Office has recently refused the copyright registration of a two-dimensional AI generated artwork entitled “A Recent Entrance to Paradise”. However, Abbott & Rothman (2022) argues for giving the copyright to AI generated works, which will encourage people to develop and use creative AI, promote transparency and eventually benefit the public interest.
4.2 Technique Related: Concept Removal, and Negative Sampling